OpenStudy (anonymous):

Simplify. Identify any x-values for which the expression is undefined. \[\frac{ x+4 }{3^{2}+11x-4 }\]

OpenStudy (anonymous):

\[\frac{ x+4 }{3^{2}+11x-4 }\]

OpenStudy (anonymous):

I think its 3x^2?

OpenStudy (anonymous):

thats not one of the options

OpenStudy (anonymous):

First factorize it...

OpenStudy (anonymous):

I mean the 3^2 in the denominator should be 3x^2

OpenStudy (anonymous):

\[\frac{ 4 }{ x+2 }\]

OpenStudy (anonymous):

\[\frac{ 3 }{ x-1 }\]

OpenStudy (anonymous):

1/(3*x-1)

OpenStudy (anonymous):

\[\frac{ 1 }{ 3x+4 }\]

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

3x^2+11x-4=(x+4)(3x-1)